# SageAttention2: Efficient Attention with Thorough Outlier Smoothing and Per-thread INT4 Quantization

### v4

@@ A.4 Experiment of Smoothing V
Table 10 shows the attention accuracy on real tensors sampled from CogvideoX with and without smoothing $V$. It demonstrates that smoothing V could improve the accuracy of SageAttention2 when quantizing $Q,K$ to INT4 and $\widetilde{P},V$ to FP8.
A.5 Per-Thread Quantization Formulation
Figure 17: Memory layout of INT4/INT8 tensor core for accumulator matrix $C$ and $D$ in $D=A*B+C$ among 32 threads (T0 $\sim$ T31) in a warp. $C$ and $D$ is of shape 16x8. Each thread only holds 4 out of the 128 elements.
@@ A.5 Per-Thread Quantization Formulation
Tensor cores, first introduced in NVIDIA’s Volta architecture, are specialized units designed for efficient matrix-multiply-and-accumulate (MMA) operations. Their usage significantly enhances computational efficiency and performance in AI and high-performance computing (HPC) workloads. Tensor cores compute small tiles of MMA operations, specifically $D=A*B+C$ on a warp (32 contiguous threads) basis. Each thread in the warp holds a fragment of input matrices and will get a fragment of output matrix as a computation result. The INT4 mma.m16n8k64 tensor core operation computes the product of a $16\times 64$ INT4 matrix $A$ and a $64\times 8$ INT4 matrix $B$, both stored in registers. It accumulates the result into a $16\times 8$ INT32 matrix $C$, also stored in registers, and returns the final product matrix $D$, which has the same shape ($16\times 8$), data type (INT32), and storage location (registers). Each thread holds only $\frac{1}{32}$ of the input and output data. Fig. 17 extracted from the PTX document (NVIDIA, ) shows the memory layout of matrix $C$ and $D$ among 32 threads in a warp. Each thread only holds 4 out of the 128 result elements.
@@
In Eq. A.5, $c_{w}$ is the count of GPU Warps. $b_{q}$ and $b_{k}$ are the block size of $Q,K$. $n$ is the token index of $Q,K$.
@@
By ensuring results held by each thread share a common dequantization scale (belong to the same quantization group), we can avoid the overhead associated with per-token quantization. Leveraging this observation, we design per-thread quantization, as shown in Fig. 4. For typical block size of $b_{q}=128$, $b_{k}=64$ and warp number $c_{w}=4$ (as used in FlashAttention2), each warp processes a tile of 32 query tokens and 64 key tokens. Query tokens $i,8+i,16+i,24+i$ ($i=0,1,\cdots,7$) can be made into one quantization group and key tokens $j,1+j,8+j,9+j,\cdots,56+j,57+j$ ($j=0,1,2,3$) can be made into one quantization group. This design aligns with the memory layout of output matrix $D$ of tensor core shown in Fig. 17, ensuring that each thread only needs one $Q$ scale and one $K$ scale for dequantization.
@@
As a result, this approach creates 32 quantization groups for $Q$ (8 for each of the 4 warps) and 4 quantization groups for $K$ in a 128x64 block, providing 32×32$\times$32 × and 4×4$\times$4 × finer granularity compared to per-block quantization for query tokens and key tokens, respectively. Table 6 and Table 15 show the accuracy gains by using per-thread quantization. Per-thread quantization achieves accuracy that closely matches per-token quantization, without introducing any kernel speed degradation (see Fig. 18).
A.6 Datasets and Metrics in Experiments
@@
Datasets. Text-to-text models are evaluated on four zero-shot tasks: WikiText (Merity et al., 2022) to assess the model’s prediction confidence, LAMBADA (Paperno et al., 2016) evaluate contextual understanding, MMLU (Hendrycks et al., 2020) for measuring knowledge across various subjects, and Longbench (Bai et al., 2024) for comprehensive assessment of long context understanding capabilities. Text-to-video models are evaluated using the open-sora (Zheng et al., 2024) prompt sets. Text-to-image models are assessed on MJHQ-30K (Li et al., 2024). TIMM is evaluated on on three image datasets: ImageNet (Deng et al., 2009), ImageNet-Sketch (Sketch) (Wang et al., 2019), and ImageNet-Rendition (ImageNet-r) (Hendrycks et al., 2021).
@@
End-to-end metrics. For text-to-text models, we use perplexity (ppl.) (Jelinek et al., 1977) for WikiText, Accuracy (Acc.) for LAMBADA and MMLU, and Longbench score (Bai et al., 2024). For text-to-video models, following Zhao et al. (2025), we evaluate the quality of generated videos on five metrics: CLIPSIM and CLIP-Temp (CLIP-T) (Liu et al., 2024) to measure the text-video alignment; (VQA-a) and (VQA-t) to assess the video aesthetic and technical quality, respectively; and Flow-score (FScore) for temporal consistency (Wu et al., 2023). For text-to-image models, generated images are compared with the images in MJHQ-30K dataset in three aspects: FID (Heusel et al., 2017) and sFID (Salimans et al., 2016) for fidelity evaluation, Clipscore (CLIP) (Hessel et al., 2021) for text-image alignment, and ImageReward (IR) (Xu et al., 2023) for human preference. For TIMM, we use classification accuracy.
@@
Smooth K (Ours) 90.86% 0.3565 0.1464
Smooth Q (Ours) 93.10% 0.2989 0.2195
@@
A.7 Additional Experiments and Analysis
@@
Results of Super-Long Context. We further conduct experiments on super-long context using Llama-3-262k (8B) [^1] https://huggingface.co/gradientai/Llama-3-8B-Instruct-262k on InfiniBench (Zhang et al., 2024) and Needle-in-a-Haystack (NIAH) (Kamradt, 2023), with sequence lengths reaching up to 262k tokens on an H100 GPU. Results are shown in Table 14 and Fig 18. SageAttention2 maintains model performance even under super-long context, while FlashAttention3(fp8) suffers from end-to-end accuracy degradation.

### v7

@@ A.4 Experiment of Smoothing V
Table 10 shows the attention accuracy on real tensors sampled from CogvideoX with and without smoothing $V$. It demonstrates that smoothing $V$ could improve the accuracy of SageAttention2 when quantizing $Q,K$ to INT4 and $\widetilde{P},V$ to FP8. We find that smoothing V is generally effective for diffusion models (Zheng et al., 2023, 2024b, 2024a, 2025; Zhao et al., 2024, 2025a; Wang et al., 2024).
A.5 Theoretical Analysis of Smoothing
In this section, we analyze the benefit of smoothing from a theoretical perspective. Let $X\in\mathbb{R}^{N\times d}$ be $N$ activation tokens of dimension $d$. Following (Dettmers et al., 2023), we suppose that an activation token follows an Gaussian distribution $\mathcal{N}(\boldsymbol{\mu},\Sigma^{2})$, where $\boldsymbol{\mu}=(\mu_{1},\mu_{2},\ldots,\mu_{d})$ and $\Sigma^{2}$ is a diagonal matrix with $\Sigma^{2}=\mathrm{diag}(\sigma_{1}^{2},\sigma_{2}^{2},\ldots,\sigma_{d}^{2})$. Further, we suppose that different token $X_{i}$ is i.i.d. sampled from the same distribution.
Suppose the absolute maximum value in a quantization group (Hu et al., 2025; Zhang et al., 2025j) is $M$, and the bit width is $b$, then there are $2^{b}$ quantization levels. Under the round-to-nearest strategy, the expected quantization error is $\frac{1}{2}\cdot\frac{2M}{2^{b}}$, which is proportional to the maximum absolute value in the quantization group. So a smaller absolute maximum value leads to smaller quantization error.
After smoothing, we have:
$Y_{ij}=X_{ij}-\frac{1}{N}\sum_{k=1}^{N}X_{kj}$ (3)
and $Y_{ij}$ also follows a Gaussian distribution. The mean and variance of $Y_{ij}$ can be calculated as follows:
$\mathbb{E}[Y_{ij}]$ $=\mathbb{E}[X_{ij}]-\frac{1}{N}\sum_{k=1}^{N}\mathbb{E}[X_{kj}]=\mu_{j}-\frac{1}{N}\sum_{k=1}^{N}\mu_{j}=0$ (4)
$\mathrm{Var}[Y_{ij}]$ $=\mathrm{Var}[\frac{N-1}{N}X_{ij}]+\sum_{k=1,k\neq i}^{N}\mathrm{Var}[\frac{1}{N}X_{kj}]=\frac{(N-1)^{2}}{N^{2}}\sigma_{j}^{2}+(N-1)\frac{1}{N^{2}}\sigma_{j}^{2}=\frac{(N-1)}{N}\sigma_{j}^{2}$ (5)
So $Y_{ij}$ has a smaller mean and variance compared to $X_{ij}$. Following the property of Gaussian distribution, we have:
$P(|Y_{ij}|<\epsilon)>P(|X_{ij}|>\epsilon),\ \forall\epsilon>0$ (6)
So the distribution of $X_{ij}$ is more concentrated toward 0 after smoothing. Then we know that
$P(\mathrm{absmax}(Y_{i})<\epsilon)=\prod_{j=1}^{d}P(|Y_{ij}|<\epsilon)>\prod_{j=1}^{d}P(|X_{ij}|>\epsilon)=P(\mathrm{absmax}(X_{i})<\epsilon)\$(7)
So this makes the distribution of absolute max value in a token more concentrated towards 0, leading to smaller quantization error.
A.6 Per-Thread Quantization Formulation
Figure 18: Memory layout of INT4/INT8 tensor core for accumulator matrix $C$ and $D$ in D=AB+CD=A*B+C among 32 threads (T0 $\sim$ T31) in a warp. $C$ and $D$ is of shape 16x8. Each thread only holds 4 out of the 128 elements.
@@ A.6 Per-Thread Quantization Formulation
Tensor cores, first introduced in NVIDIA’s Volta architecture, are specialized units designed for efficient matrix-multiply-and-accumulate (MMA) operations. Their usage significantly enhances computational efficiency and performance in AI and high-performance computing (HPC) workloads. Tensor cores compute small tiles of MMA operations, specifically D=AB+CD=A*B+C on a warp (32 contiguous threads) basis. Each thread in the warp holds a fragment of input matrices and will get a fragment of output matrix as a computation result. The INT4 mma.m16n8k64 tensor core operation computes the product of a $16\times 64$ INT4 matrix $A$ and a $64\times 8$ INT4 matrix $B$, both stored in registers. It accumulates the result into a $16\times 8$ INT32 matrix $C$, also stored in registers, and returns the final product matrix $D$, which has the same shape ($16\times 8$), data type (INT32), and storage location (registers). Each thread holds only $\frac{1}{32}$ of the input and output data. Fig. 18 extracted from the PTX document (NVIDIA, ) shows the memory layout of matrix $C$ and $D$ among 32 threads in a warp. Each thread only holds 4 out of the 128 result elements.
@@
By ensuring results held by each thread share a common dequantization scale (belong to the same quantization group), we can avoid the overhead associated with per-token quantization. Leveraging this observation, we design per-thread quantization as formulated in Eq. A.6, where $c_{w}$ is the count of GPU Warps, $b_{q}$ and $b_{k}$ are the block size of $Q,K$, and $n$ is the token index of $Q,K$. For typical block size of $b_{q}=128$, $b_{k}=64$ and warp number $c_{w}=4$ (as used in FlashAttention2), each warp processes a tile of 32 query tokens and 64 key tokens. Query tokens $i,8+i,16+i,24+i$ ($i=0,1,\cdots,7$) can be made into one quantization group and key tokens $j,1+j,8+j,9+j,\cdots,56+j,57+j$ ($j=0,1,2,3$) can be made into one quantization group, as visualized in Fig. 4. This design aligns with the memory layout of output matrix $D$ of tensor core shown in Fig. 18, ensuring that each thread only needs one $Q$ scale and one $K$ scale for dequantization.
@@
As a result, this approach creates 32 quantization groups for $Q$ (8 for each of the 4 warps) and 4 quantization groups for $K$ in a 128x64 block, providing $32\times$ and $4\times$ finer granularity compared to per-block quantization for query tokens and key tokens, respectively. Table 6 and Table 15 show the accuracy gains by using per-thread quantization. Per-thread quantization achieves accuracy that closely matches per-token quantization, without introducing any kernel speed degradation (see Table 19 and 19).
A.7 Datasets and Metrics in Experiments
@@
Datasets. Text-to-text models are evaluated on four zero-shot tasks: WikiText (Merity et al., 2022) to assess the model’s prediction confidence, LAMBADA (Paperno et al., 2016) evaluate contextual understanding, MMLU (Hendrycks et al., 2021b) for measuring knowledge across various subjects, and Longbench (Bai et al., 2024) for comprehensive assessment of long context understanding capabilities. Text-to-video models are evaluated using the open-sora (Zheng et al., 2024c) prompt sets. Text-to-image models are assessed on MJHQ-30K (Li et al., 2024). TIMM is evaluated on on three image datasets: ImageNet (Deng et al., 2009), ImageNet-Sketch (Sketch) (Wang et al., 2019), and ImageNet-Rendition (ImageNet-r) (Hendrycks et al., 2021a).
@@
End-to-end metrics. For text-to-text models, we use perplexity (ppl.) (Jelinek et al., 1977) for WikiText, Accuracy (Acc.) for LAMBADA and MMLU, and Longbench score (Bai et al., 2024). For text-to-video models, following Zhao et al. (2025b), we evaluate the quality of generated videos on five metrics: CLIPSIM and CLIP-Temp (CLIP-T) (Liu et al., 2024) to measure the text-video alignment; (VQA-a) and (VQA-t) to assess the video aesthetic and technical quality, respectively; and Flow-score (FScore) for temporal consistency (Wu et al., 2023). For text-to-image models, generated images are compared with the images in MJHQ-30K dataset in three aspects: FID (Heusel et al., 2017) and sFID (Salimans et al., 2016) for fidelity evaluation, Clipscore (CLIP) (Hessel et al., 2021) for text-image alignment, and ImageReward (IR) (Xu et al., 2023) for human preference. For TIMM, we use classification accuracy.
@@
A.8 Kernel Benchmark Setup
We benchmark kernel speed with a batch size of 4 and 32 attention heads across a variety of sequence lengths. Benchmarks are conducted using head dimensions of 64 and 128, both with and without Causal Mask (Vaswani, 2017). To generate input tensors for benchmarking, we follow standard practices adopted in prior works such as FlashAttention (Dao et al., 2022). For floating-point data types, inputs are drawn from a Gaussian distribution with mean 0 and standard deviation 1, while for integer data types, inputs are uniformly sampled within the representation range:[-128, 127] for INT8 and [-8, 7] for INT4.
@@ A.8 Kernel Benchmark Setup
Smooth K 90.86% 0.3565 0.1464
Smooth Q 93.10% 0.2989 0.2195
@@
Table 19: Comparison of different quantization granularities measured on L20 GPU, with $QK^{\top}$ in INT4 and $\widetilde{P}V$ in FP8.
   Granularity    TOPS
   Per-tensor    286
   Per-block    284
   Per-thread    283
   Per-token    268
A.9 Additional Experiments and Analysis
@@
Results of Super-Long Context. We further conduct experiments on super-long context using Llama-3-262k (8B) [^1] https://huggingface.co/gradientai/Llama-3-8B-Instruct-262k on InfiniBench (Zhang et al., 2024) and Needle-in-a-Haystack (NIAH) (Kamradt, 2023), with sequence lengths reaching up to 262k tokens on an H100 GPU. Since Hopper GPUs lack native INT4 tensor core support, we use SageAttention2-8b for this evaluation. We compare it against FlashAttention3(fp8), ensuring both methods operate under the same bit width. Results are shown in Table 14 and Fig 19. SageAttention2 maintains model performance even under super-long context, while FlashAttention3(fp8) suffers from end-to-end accuracy degradation.
Results of Audio Tasks. We evaluate Qwen2-Audio (7b) (Chu et al., 2024), a speech-to-text model, on the ASR task using the Librispeech (Panayotov et al., 2015) test split and measured its performance with the WER metric (Word Error Rate). As shown in Table 20, SageAttention2 consistently outperforms the baselines, highlighting its effectiveness in audio-related models and benchmarks.
Table 20: End-to-end metrics on Qwen2-Audio (7B).
Model Attention Test-Clean $\downarrow$ Test-Dev $\downarrow$
Qwen2-Audio Full-Precision 1.74 4.01
HadmdAttn 1.77 4.05
SmoothAttn 1.76 4.01
SageAttention 1.74 4.02
SageAttn2-4b 1.73 3.99
SageAttn2-8b 1.72 4.03